OpenStudy (anonymous):

if f(x)=(x+3)^2 then f(x+h)-f(x)/h = Ax+Bh+C Find constants A, B, and C.

OpenStudy (anonymous):

when i plug the first thing into the second everything cancels out

OpenStudy (sasogeek):

do you know about differentiation by partial fractions?

OpenStudy (anonymous):

dont think so. i just know how to do rate of change and how to take the derivative of something...

myininaya (myininaya):

what is \[f(x+h)\] given that \[f(x)=(x+3)^2\]

OpenStudy (anonymous):

((x+3)^2) + h) ???

myininaya (myininaya):

you just replace x with x+h

OpenStudy (sasogeek):

what that means is that, (x+h) imply x, so where ever u see x, replace it with (x+h) :)

OpenStudy (amistre64):

i think they should not use x in both cases; when x = a+h might be less confusing

OpenStudy (sasogeek):

true

OpenStudy (anonymous):

i think thats what i did and everything canceled out on the left side

OpenStudy (amistre64):

if f(x)=(x+3)^2 then when f(a) and f(a+h)

myininaya (myininaya):

amistre you want a function of x not a since both sides is in terms of x hehe :) you can use your fancy a if you want to though

OpenStudy (amistre64):

x = a ;)

OpenStudy (amistre64):

we use substitution all the time in calculus and diffies

OpenStudy (sasogeek):

\[\large \frac{(x+h+3)^2-(x+3)^2}{h} \]

OpenStudy (anonymous):

hmm i see

myininaya (myininaya):

now plug multiply that top junk out and some stuff will zero out

OpenStudy (sasogeek):

\[\large \frac{(x+h+3)^2-(x+3)^2}{h}=Ax+Bh+C\]

myininaya (myininaya):

then you will be able to factor out in h on top

myininaya (myininaya):

then you will be like oh h/h=1 i don't have that crazy h on bottom anymore

OpenStudy (anonymous):

so i would get 1 on the left?

OpenStudy (anonymous):

if so i did it differently and still got 1

myininaya (myininaya):

\[\frac{(x+3+h)^2-(x+3)^2}{h}=\frac{(x+3)^2+2(x+3)h+h^2-(x+3)^2}{h}\]

myininaya (myininaya):

\[=\frac{(x+3)^2-(x+3)^2+2h(x+3)+h^2}{h}\]

myininaya (myininaya):

\[=\frac{0+2h(x+3)+h^2}{h}=\frac{2h(x+3)+h^2}{h}=\frac{h(2(x+3)+h)}{h}\] do you see what i mean about the h/h thing now?

OpenStudy (anonymous):

yeah...

OpenStudy (sasogeek):

you'll have so many of the h's cancelling out, and the rest will be easy to figure out, equate your final equations to the right hand side, I hated partial fractions so i'll probably have trouble with it though

OpenStudy (anonymous):

could you tell me how you foiled this (x+h+3)^2 or whatever?

OpenStudy (sasogeek):

what do you mean by foil... the writing or how that expression came about?

OpenStudy (anonymous):

or simplify...

OpenStudy (anonymous):

if there was only two things in the paranthesis, you woul foil, but what would you do with this

OpenStudy (asnaseer):

you could take advantage of the fact that you have the difference of two squares in the expression to simplify first.

myininaya (myininaya):

\[(a+b)^2=a^2+2ab+b^2\] \[(x+3+h)^2=(x+3)^2+2(x+3)h+h^2\]

OpenStudy (sasogeek):

\[\huge (a+b)^2=a^2+2ab+b^2 \]\[\huge a=x+3 \]\[\huge b=h \]

OpenStudy (asnaseer):

\[\frac{(x+3+h)^2-(x+3)^2}{h}=\frac{(x+3+h+x+3)(x+3+h-x-3)}{h}\]\[=\frac{(2x+6+h)h}{h}=2x+6+h\]

OpenStudy (asnaseer):

this uses:\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (sasogeek):

Naseer I believe she was referring to " (x+h+3)^2 " which she stated lol :P but yeah what u said is true though, i didn't even notice it myself :P good to know :D

OpenStudy (asnaseer):

yes saso - I was just highlighting a simpler way of getting to the same answer that both you and myininaya got to

myininaya (myininaya):

whatever simpler my answer rules!

OpenStudy (anonymous):

me = he

OpenStudy (sasogeek):

always taking the shortcuts huh :P

OpenStudy (asnaseer):

:)

OpenStudy (sasogeek):

my bad bro, sorry lol

OpenStudy (anonymous):

and im still a little lost but if stare at this a little longer ill get it down in my head lol

OpenStudy (sasogeek):

sure :)

OpenStudy (asnaseer):

I can redo the steps from the beginning for you if it helps mario?

OpenStudy (asnaseer):

I know it can be confusing when you get lots of ideas thrown at you

OpenStudy (anonymous):

if it'll make me understand everything better.

OpenStudy (anonymous):

we just started doing all of this in class so its new material to me

OpenStudy (sasogeek):

i love this guys enthusiasm to study lol, keep it up mario :D

OpenStudy (asnaseer):

ok, we started with:\[f(x)=(x+3)^2\]using this we get:\[f(x+h)=(x+h+3)^2\]

OpenStudy (asnaseer):

therefore:\[f(x+h)-f(x)=(x+h+3)^2-(x+3)^2\]\[=(x+h+3+x+3)(x+h+3-x-3)\]\[=(2x+h+6)h\]

OpenStudy (asnaseer):

all ok so far?

OpenStudy (anonymous):

=(x+h+3+x+3)(x+h+3−x−3) =(2x+h+6)h dont understand this stuff

OpenStudy (anonymous):

what did you do to get from point a to point b

OpenStudy (asnaseer):

there I used the fact that:\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (asnaseer):

where a in this case is x+h+3 and b is x+3

OpenStudy (asnaseer):

are you familiar with this rule?

OpenStudy (anonymous):

a little, yes

OpenStudy (anonymous):

how did you know which one is a and which one is b

OpenStudy (asnaseer):

\[a^2-b^2\]is compared to:\[(x+h+3)^2-(x+3)^2\]

OpenStudy (asnaseer):

so \(a^2\) is the term on the left and \(b^2\) is the term on the right

OpenStudy (anonymous):

ok i see i see

OpenStudy (asnaseer):

ok, so we got to:\[f(x+h)-f(x)==(2x+h+6)h\]divide both sides by h to get:\[\frac{f(x+h)-f(x)}{h}=2x+h+6\]

OpenStudy (asnaseer):

you are told that this also equals:\[Ax+Bh+C\]therefore:\[2x+h+6=Ax+Bh+C\]now just compare the coefficients of x and h to determine A and B and then C is just the constant term.

OpenStudy (anonymous):

that was a really good explanation. thanks a lot!

OpenStudy (asnaseer):

yw

OpenStudy (asnaseer):

and sorry to saso and myininaya for stepping on your shoes :(

OpenStudy (anonymous):

haha

myininaya (myininaya):

i still like my way better :)

OpenStudy (asnaseer):

I will let the lady take the high ground :)

OpenStudy (anonymous):

sasogeek, myininaya, and amistre, thank you as well :)